OpenStudy (anonymous):

Medal given if mobile allows me to To sum it up: we have to use the formula h = v^2 over 64. The height is 16.5 How fast was the record holder running? Round to the nearest tenth.

OpenStudy (texaschic101):

h = v^2 / 64 h = 16.5 now we sub 16.5 = v^2/64 --- multiply both sides by 64 1056 = v^2 -- now take the square root of both sides sq rt 1056 = v -- notice by taking the sq rt of v^2, it removes the ^2 32.496 (rounds to 32.5) = v check.. 16.5 = 32.5^2 / 64 16.5 = 1056.25 / 64 16.5 = 16.5 (correct) so, v = 32.5